OpenStudy (anonymous):

Find f(-3). f(x) = 4x + 7 A) -5 B) 6 C) 10 D) 19

OpenStudy (anonymous):

WOW. What's up with the spacing?

OpenStudy (anonymous):

just plug the -3 into the x

OpenStudy (anonymous):

i dont know why it done tht

OpenStudy (anonymous):

so you get 4(-3)+7

OpenStudy (anonymous):

Replace the x with -3 4(-3)+7= your answer.

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

What is -4*3=.........Now add 7 to that answer.

OpenStudy (anonymous):

19

OpenStudy (anonymous):

-12+7=?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

It will give you a negative number.

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

If r = -6, s = -2, t = -3 v = 10, and w = - 1 2 then wv|-t - s| = A) -25 B) -5 C) 5 D) 25